OpenStudy (anonymous):

what mean z^2.. it mean square like usual? this topic in complex number

OpenStudy (jhannybean):

I'm not understanding your question...

jimthompson5910 (jim_thompson5910):

if z = a+bi then z^2 = (a+bi)^2 = a^2 + 2abi + b^2i^2 but remember that i^2 = -1, so b^2i^2 turns into b^2(-1) = -b^2

jimthompson5910 (jim_thompson5910):

so z^2 = a^2 + 2abi - b^2

OpenStudy (anonymous):

thnx

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

if this quetion want values of z and it give z^2= 8+6i??

jimthompson5910 (jim_thompson5910):

this would mean that z^2= 8+6i a^2 + 2abi - b^2 = 8+6i separate the real and imaginary parts to get a^2 - b^2 = 8 2abi = 6i ---> 2ab = 6

jimthompson5910 (jim_thompson5910):

so you have a system of equations, which you can use to solve for a and b

jimthompson5910 (jim_thompson5910):

effectively, you are finding the square root of w = 8+6i ie, z^2 = w